What is the length of a flag with a 15 inch diagonal and 9 inch height?
( I just started learning Pythagorean theorem and I realize the formula is a^2 + b^2 = c^2. Would that mean C^2 is 15^2 and a^2 is 9^2? Then you basically just try to find A by subtracting C and A then figuring out what number times itself equals the answer? )

Answer by jim_thompson5910(35256)

You have the right idea. However, we are looking for 'b' if you made a = 9.


a^2 + b^2 = c^2

9^2 + b^2 = 15^2

81 + b^2 = 225

b^2 = 225-81 ... Subtract 81 from both sides.

b^2 = 144

b = sqrt(144) ... Take the square root of both sides (note: b > 0)

b = 12


So the length is 12 inches